OpenStudy (anonymous):

What is the range of f(x)=√(3x-7) -0.5? I got [3,∞) which I believe is correct for the domain, but I'm not sure what the range would be and if it would be [-0.5,∞) or not

OpenStudy (anonymous):

Or if it's [0.91, ∞)?

OpenStudy (jdoe0001):

\(\bf f(x) = \sqrt{ (3x-7) } - 0.5 \quad ?\)

OpenStudy (anonymous):

Yes that's the equation

OpenStudy (jdoe0001):

well, with EVEN radicals, that is, like square root, or root 2 <-- even is that you cannot have a negative value inside it, otherwise you'd end up with an "imaginary" value so the value in it has to be positive so any values "x" can take, MUST NOT make the radical quantity, 3x - 7, negative 0 is ok, -1 or less is not good, because \(\bf \sqrt{ 0 } = 0\) , so that's ok so at what point will that be for "x"? well, we can easily find out by setting the equation to 0 3x - 7 = 0 3x = 7 \(\bf x = \frac{ 3 }{ 7 }\) so, the lowest "x" can go is \(\bf \frac{ 3 }{ 7 }\) , at that point, the radical gives 0 and y = -0.5 so "x" cannot go lower than \(\bf \frac{ 3 }{ 7 }\) , it can go higher though, up to \(\bf +\infty\)

OpenStudy (jdoe0001):

ohh... I messed one thing.... \(\bf x = \frac{ 7 }{ 3 }\) rather

OpenStudy (anonymous):

So the range would then be [7/3, ∞)?

OpenStudy (jdoe0001):

yeap

OpenStudy (anonymous):

Thank you!

OpenStudy (jdoe0001):

yw

OpenStudy (jdoe0001):

at that point, y = -0.5 from there, as the radical gives a positive value, then "y" goes upwards

OpenStudy (jdoe0001):

well, a plus version and minus version will come out of the radical...so the range will ... go upwards and downwards

OpenStudy (anonymous):

So, the domain is [3, ∞) and the range is [7/3, ∞). I think that sounds right

OpenStudy (jdoe0001):

So the range would then be [7/3, ∞)? <--- ... I misread you.... that's correct for the DOMAIN

OpenStudy (anonymous):

Ohhh.... okay, sorry, it got confusing

OpenStudy (jdoe0001):

the values "x" takes, are the domain of the function

OpenStudy (anonymous):

and [-0.5, ∞) is the range, thusly? Is that correct?

OpenStudy (jdoe0001):

recall that \(\bf f(x) = \sqrt{ (3x-7) } - 0.5 \implies f(x) = \pm \sqrt{ (3x-7) } - 0.5\) so you see, there's a + and a - version from the radical so I'd say it goes to \(\boldsymbol{(-\infty, +\infty)}\)

OpenStudy (anonymous):

But when x is 2 and below, there is no y value, so wouldn't that make the domain never be negative?

OpenStudy (jdoe0001):

I'm afraid not, the EVEN roots, have 2 values from it, a + and a -, so any value resulting from a x = 5, 25, 1000.... will yield a +y and -y exact value keep also in mind that \(\bf f(x) = \sqrt{ (3x-7) } - 0.5 \implies f(x) = \pm \sqrt{ (3x-7) } - 0.5\\ \quad \\ \quad \\ y = \sqrt{ (3x-7) } - \cfrac{ 1 }{ 2 } \implies y + \cfrac{ 1 }{ 2 }= \sqrt{ (3x-7) }\\ \implies \left(y + \cfrac{ 1 }{ 2 }\right)^2=3x-7\\\quad \\ \cfrac{ 1 }{ 3 }\left(y + \cfrac{ 1 }{ 2 }\right)^2+\cfrac{ 7 }{ 3 }=x\) is a parabola

OpenStudy (anonymous):

I'm a still a little confused but alright

OpenStudy (jdoe0001):

notice the 2 + and - roots

OpenStudy (anonymous):

I only need the positive graph of it, though. Not the negative square root graph.

OpenStudy (jdoe0001):

hmm... well... , I'd think you need both.... but maybe the exercise is only concerned with the positive one

OpenStudy (anonymous):

yeah, it is only concerned with the positive one

OpenStudy (jdoe0001):

ok... then your interval of \(\boldsymbol{ [-0.5, +\infty) }\) is correct

OpenStudy (anonymous):

Oh, ok, good, thanks for all the help!

OpenStudy (jdoe0001):

yw